The numerical values of the three sides of a triangle are given by three consecutive odd integers. If the perimeter is greater than or equal to 15 inches, what are the possibilities for the shortest side?
The length of a rectangle is 4 times the width. If the perimeter is to be greater than or equal to 170 meters. What are the possible values for the width?
PLEASE HELP ME :(

Answer by Alan3354(69443)

The numerical values of the three sides of a triangle are given by three consecutive odd integers. If the perimeter is greater than or equal to 15 inches, what are the possibilities for the shortest side?
--------------
3 + 5 + 7 = 15
Short side >= 3
--------
The length of a rectangle is 4 times the width. If the perimeter is to be greater than or equal to 170 meters. What are the possible values for the width?
----
L = 4W
2L + 2W >= 170
L + W >= 85
---
4W + W >= 85
W >= 17